Fermi National Laboratory
 

Accelerator Update for Jun 12, 2006 - June 13, 2006

 

Monday June 12

The day shift began with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), Recycler, and Tevatron (TeV) conducting studies, with experts investigating the NuMI Horn power supply ground fault, and with MiniBooNE taking beam.

At 10:43 AM, the TeV suffered a sector A1 quench during an intentional abort of protons.

Operations reported at 10:59 AM that the NuMI Horn hadn’t tripped off for a while. A NuMI expert asked for some low intensity beam.

At 11:14 AM, Operations reported that the NuMI beam permit was being held off on a target and water skid OK bit and an E:V118PS regulation bit. Operations contacted experts.

Operations sent beam to NuMI at 11:50 AM.

Operations prepared Meson enclosures M01 – M03 for an access at 12:35 PM.

Operations put the TeV into a dry squeeze at 12:44 PM.

Recycler experts began a study period around 1 PM.

Operations began preparing the Recycler at 4:28 PM to stash antiprotons.

RF experts completed changing the phase detector for two MI RF stations (MIRF8 & 11) at 5:12 PM.

At 5:57 PM, the 400MeV Klystron Debuncher cavity water skid tripped off on a LCW flow fault. Operators turned the pump back on.

The Recycler stash ended around 6 PM.

Operations put the TeV into shot setup around 7:19 PM.

Operations established store 4759 at 11:46 PM with an initial luminosity of 25.4E30. Pbar resumed stacking.

Tuesday June 13

The midnight shift began with Operations monitoring store 4759 and a small stash, with Pbar stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with MiniBooNE and NuMI taking beam.

At 12:47 AM, the Klystron Debuncher cavity water skid tripped off on a water flow fault for the first of twenty times in a little over five hours. It reset each time.

Recycler began preparing to stash at 4:45 AM.

The stash ended and Pbar resumed stacking at 6:29 AM.

The Plans for Tuesday

The plans for today are to leave the store in until later this afternoon, and to try and resolve the sequencer and SDA problems. Booster experts will be allowed to conduct some destructive orbit and extraction tuning. Pbar experts will be allowed to conduct target spot size measurements.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/13/2006    email Fermilab